# Individual or collective treatments: how to target antimicrobial use to limit the spread of Mannheimia haemolytica among beef cattle?

### v1

@@ Results
Every scenario using a collective treatment yielded a lower cumulative incidence of BRD cases than their counterpart scenario using individual treatment only, while displaying as expected larger AMU (Figure 1). The difference in cumulative incidence of BRD cases between using individual treatments and collective treatments was however lower in HR30 than in HR90. In HR30, each collective treatment displayed roughly an equal median cumulative incidence of BRD cases. However, the one triggered by the slope criterion had 18 to 33% lower AMU than the other two criteria. In HR90, the scenario using the slope criterion and the conventional collective treatment displayed a lower median cumulative incidence of BRD cases than the severity criterion. However, the scenario with the slope-based criterion yielded a 10% lower AMU than the conventional collective treatment for a roughly equivalent cumulative incidence. Scenarios with sorted pens tended to yield smaller AMU than their random counterparts.
@@
Figure 1: Distributions of the antimicrobial use (AMU) vs cumulative incidence of BRD cases in the building population. The darker the shade, the higher the density of replicates bearing this output combination. The crosses represent the median of the distributions. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : the population is composed of 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: the population is composed of 10% of low risk animals and 90% of high risk animals
The treatments appeared to have only a limited impact on the average duration of severe clinical signs (Figure 2). Indeed, in most scenarios, we observed an important overlap of the distributions on the y-axis.
Figure 2: Distributions of the antimicrobial use (AMU) vs average duration with severe clinical signs for 200 animals. The darker the shade, the higher the density of replicates bearing this output combination. The crosses represent the median of the distributions. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : the population is composed of 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: the population is composed of 10% of low risk animals and 90% of high risk animals
The treatment false discovery rate (FDR) was very low when only individual treatments were used and way higher for all collective treatments (Figure 3). In HR30, the median FDR of collective treatments was between 50 and 75%. The collective treatment with the lowest FDR appeared to be the one triggered by the slope criterion. Sorting pens did not appear to have a major effect on treatment FDR. Conversely, with HR90, the conventional collective treatment had the highest FDR and the one triggered by the severity criterion had the lowest.
Figure 3: Distributions of the treatment False Discovery Rate (FDR). The FDR is the ratio of doses given to healthy individuals on the total AMU. The dashed lines are the median of the distributions. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : the population is composed of 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: the population is composed of 10% of low risk animals and 90% of high risk animals.
The criterion for triggering a collective treatment affected the delay between the detection of the first BRD case and the collective treatment (Figure 4). Indeed, when collective treatments where triggered with the slope criterion, the delay between the first case and the collective treatment was shorter than a day in the majority of the stochastic replicates, whereas this delay was more often between one and two days for the conventional criterion. Collective treatment triggered by the severity criterion were triggered in less than a day between 26% and 37% of the replicates with collective treatment. In around 43% of the cases, it was triggered between one and two days after the first case in the pen.
Figure 4: Delay between the beginning of an episode and the collective treatment. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row: the population is composed of 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: the population is composed of 10% of low risk animals and 90% of high risk animals.
On an individual level, collective treatments shortened the delay between the infection and the treatment 5. The delay exceeded 2 days in less than 20% of the replicates in scenarios with collective treatments, while delays up to 9 days were observed in scenarios with individual treatments. Among collective treatment criteria, the conventional criterion yielded the biggest proportions of delays shorter than a day (up to 52 % in HR90 scenarios), while the other two criteria yielded similar results.
@@
Figure 5: Delay between infection and treatment. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row: the population is composed of 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: the population is composed of 10% of low risk animals and 90% of high risk animals.
@@
Pens containing animals of the same individual risk level were grouped together across the simulations to observe proportions of triggered collective treatment in each specific group (Table 1). A collective treatment was triggered in the most of the pens composed of animals with high risk of respiratory disease, irrespective of the criterion. The conventional criterion had the highest proportion of triggered collective treatments in pens composed of medium risk individuals (87%), while the severity criterion triggered a collective treatment in 58% of these pens and the slope criterion only triggered a collective treatment in 33% of them. In pens composed of low risk individuals, collective treatments triggered by the conventional criterion happened in 37% of the pens, while collective treatments based on the severity and slope criteria only happened in 14 and 4% of the pens respectively.
@@
Table 1: Proportion of triggered collective treatments in sorted batches.
@@ Discussion
This study aimed at modeling the impact of triggering collective antimicrobial treatments according to three different criteria on the spread of M. haemolytica in multi-pens fattening systems while accounting for the multi-pen structure, individual risk of being affected by BRD and farming practices in terms of pens composition. We computed the cumulative incidence and the average duration of severe clinical signs. We put these outputs in balance with the AMU and treatment precision for each scenario. We also computed the delay between the beginning of an episode and the beginning of the treatment and calculated the proportion of pens triggering collective treatments based on each criterion. We proposed that a novel criterion, based on the speed of the spread rather than on a proportion of detected individuals, could yield a better trade-off between BRD cases incidence and AMU than the conventional way of implementing collective treatment. This result could have interesting operational implications on the decision-making regarding collective treatment in fattening farms.
@@
However, the supplementary cost in antimicrobial was less desirable when the general risk was lower in the population, as the reduction in incidence and severity thanks to collective treatment was less substantial. This suggests that a characterization of the risk level of the population could be beneficial for making the best decision regarding treatment and herd management ([edwards_control_2010]).
@@
Our model allowed us to identify a beneficial novel criterion for collective treatment. Indeed, instead of a threshold of detected animals in a pen, basing a collective treatment decision on the speed of the spread of the disease has proved to reduce the incidence and AMU. A field meta-analysis of randomized clinical trials showed that the risk reduction of BRD varied drastically with the criterion for collective treatment ([baptiste_antimicrobial_2017]). The routine criterion is 10% of morbidity for two consecutive days, however, there is no evidence based studies proving the relevance of this threshold ([radostits_herd_2001]). Perspectives of this study include the investigation of the novel treatment strategies in broader contexts in order to translate the thresholds used into operative indicators.
@@
Collective treatments based on the speed of the spread of the pathogen proved to be triggered sooner than with the other criteria. This precocity was key for the reduction of the incidence and severity of the disease, especially in high-risk populations. It also showed to target more specifically high-risk groups than the other criteria, as it reduced the proportion of low-risk pens having received collective treatments. On an individual level, the delay between the infection and the administration of the first dose of treatment was also characterized. This delay was shortened by the implementation of collective treatment strategies, particularly for the conventional criterion. These are original findings with clear practical implications, especially with regards to the need for a better knowledge by farmers of individual risk of animals developing BRD, in order to account for this information when constituting pens.
@@
Our study advocates for more transparency on the individual risk level for BRD in the cattle production system. Such a characterization could allow mindful investments for the farmers as well as more adequate control measures regarding housing and treatment. Indeed, our novel collective treatment criterion proved to be able to target at-risk pens while not applying treatments to lower risk pens.
@@ Acknowledgements
The authors declare that they comply with the PCI rule of having no financial conflicts of interest in relation to the content of the article.
@@
The complete code for the model model can be found on this git repository: https://forgemia.inra.fr/dynamo/brd-models/brd-public/-/tree/PCI_animal_science2024. The version of the model used for this work is tagged under pci_animal_science_2024.
Annexe A SI1: Model graphic overview
This supplementary information provides graphic overview of the states and processes of the model. We propose a formalism featuring the transitions from one state to another as well as the regulatory links existing between transitions. In this formalism, full arrows represent a transition from one state to another whereas the dashed lines represent regulatory links. Let $X$ be a regulator (state or transition) of transition $T_{2}$. This will be represented by a dashed arrow starting from $X$ and targeting $T_{2}$.
A.1 Model overview for M.haemolytica
Figure 6: Graphic overview of the states and transitions in each process of the model. Transitions are represented by full arrows, while dashed lines represent regulatory links, from [sorin-dupont_modeling_2023]
Annexe B SI2: Model processes
B.1 Processes of the model
@@
In the model, five processes drive the states of individuals: infection, hyperthermia, clinical signs, detection and treatment. They are represented by a formalism broadly used in computer science, finite state machines, which is close to flow diagrams used by epidemiologists, with a higher expressiveness. Their textual description in the YAML format in the model file is easy to read and can be automatically converted in a graphical representation by EMULSION. The model can be found on a git repository, accessible by following this link: https://forgemia.inra.fr/spicault/brd-withinherd-emulsion.git. The version of the model used for this work is tagged under pci_animal_science_2024.
@@
To help understand the state machine diagrams, we present how the diagram summarising the infection process relates to the state machine specification, extracted from the whole model. More details on how state machines differ from flow diagrams are available in the documentation: https://sourcesup.renater.fr/www/emulsion-public/pages/Modelling_principles.html. The following subsections are directly drawn from Sorin-Dupont et al. 2023 [sorin-dupont_modeling_2023] as they describe processes and transitions that remained unchanged from this previous version of the model.
B.2 Hyperthermia
Hyperthermia was composed of two states: hyperthermic (H) and non-hyperthermic (NH). NH animals could transition to H with probability $p_{H}$ attributed to non-infectious factors. Once in state H, they remained there for a period $\tau_{H}$ sampled from a Beta distribution adjusted using observed data, before reverting to NH. Additionally, the transitions from NH to H and back could result from the infection process.
B.3 Infection and clinical signs status
Four health statuses were considered: susceptible (S), asymptomatic carrier (E), infectious (I) and resistant (R) animals. Asymptomatic carriers could spontaneously turn I with probability $p_{E}$ and could also be infected by surrounding infectious individuals (I). Three actions were triggered when entering the state I: (1) the individual exhibited mild clinical signs for a duration $\tau_{M}$ drawn from a Beta distribution calibrated from observed data, (2) the animal changed from NH to H state, (3) a random draw with probability $p_{C}$ drove whether the individual would display severe clinical signs at the end of its mild clinical signs. If displaying severe clinical signs, a boolean deciding on the survival of the individual was drawn from a binomial law of probability $p_{d}$. Death occurred at the end of the severe clinical signs duration ($\tau_{C}$). If the individual did not die from infection, it then recovered and became resistant (R). Recovery occurred after duration $\tau_{I}$ drawn from a gamma distribution according to the given bacterial pathogen. Theoretically, $\tau_{I}$ is longer than $\tau_{M}$. However, $\tau_{M}+\tau_{C}$ could exceed $\tau_{I}$. In that case, the infectious period was $\tau_{M}+\tau_{C}$. When transitioning to R state, animals changed from H to NH.
B.4 Detection status
Two detection methods were used. The first detection relied on visual on-farm appraisal of clinical signs, assuming lethargy was the most significant sign to calibrate the delay ($\tau_{M}$) between infection and severe sign occurrence. Severe clinical signs were detected with a sensitivity of 1, while the sensitivity for mild clinical signs detection was assumed to be 0.5. The model assumed a clinical check-up at every time step (12 hours). Following the detection of the first case through visual appraisal, all hyperthermic animals were identified and transitioned frome undetected state to detected using rectal temperature measured at the next feeding time, 12 hours later.
B.5 Treatment status
Each animal detected as diseased or hyperthermic, or in the context of collective treatment transitioned from not treated (NT) to treated (T). Treated animals received one antibiotic dose, assumed to be effective after a certain duration $\tau_{T}$. If animals still exhibited clinical signs after this duration, they would be treated again for the same duration, but the number of treatments per individual per episode was limited ($max_{T}$). Transitions from T to NT occurred in three cases: (1) recovery after $\tau_{T}$ due to successful treatment with probability $p_{T}$, (2) the end of the infectious period occurred while under treatment but was not caused by it, (3) treatment failure after $max_{T}$ doses. A treatment success triggered the transition from I to R, consequently trigerring transitions driven by the end of the infectious state i.e. the transitions to non-hyperthermic and to asymptomatic states.
Annexe C SI3: Threshold selection for the Novel criteria for collective treatment
C.1 Method

### v2

@@ Results
Every scenario using a collective treatment yielded a lower cumulative incidence of BRD cases than their counterpart scenario using individual treatment only (Wilcoxon test, p-value < 0.05). The difference in cumulative incidence of BRD cases between using individual treatments and collective treatments was however lower in HR30 than in HR90 (Figure 2).
In HR30, the cumulative incidence was the lowest in scenarios applying the conventional collective treatment. In the scenario with a random pen allocation, the severity criterion had a lower median incidence than the slope criterion, whereas no significant difference between the two new criteria could be assessed in the scenario with sorted pens.
In HR90, the scenario using the slope criterion and the conventional collective treatment displayed a lower median cumulative incidence of BRD cases than the severity criterion. In the scenario with random pen allocation, no significant difference could be assessed between the conventional collective treatment criterion and the slope-based criterion whereas the latter had a significantly lower median when the pens were sorted by risk levels. Overall, sorting the pens significantly reduced the cumulative incidence in every scenario, except in HR90 for the severity criterion (Wilcoxon test, p-value <0.05).
@@
Figure 2: Distributions of the cumulative incidence of BRD cases in the building population. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
Conversely, collective treatment yielded large AMU than individual treatments (Figure 3). In every scenario, the slope criterion yielded the lowest AMU of all the collective treatment criteria, while the conventional criterion yielded the highest.
@@
Figure 3: Distributions of the antimicrobial use (AMU) in the building population. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
The treatments appeared to have only a limited impact on the average duration of severe clinical signs (Figure 4). Indeed, in most scenarios, we observed an important overlap of the distributions. However, the conventional criterion for collective treatment was generally associated with the lowest average duration of severe clinical signs.
Figure 4: Distributions of the average duration of severe clinical signs in the building population. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
Collective treatments successfully reduced the mortality in HR90 scenarios when compared to individual treatments (Figure 5). The effect was however more contrasted in HR30 scenarios, with only the mortality associated with the conventional collective treatment being significantly lower than the individual treatment in both Sorted and Random scenarios (p-value <0.05). The statistical comparisons within collective treatment strategies yielded unconclusive results with regards to significance.
Figure 5: Boxplots of the mortality in the building population. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
The proportion of treatment misuse was very low when only individual treatments were used and much higher for all collective treatments (Kruskal-Wallis test, p-value <0.05) (Figure 6). In HR30, the median misuse of antimicrobial for collective treatments was between 50 and 75%. The collective treatment with the lowest misuse appeared to be the one triggered by the slope criterion. Sorting pens appeared to have an effect on treatment misuse when collective treatments were used (Wilcoxon test, p-value <0.05). In HR90 scenarios, the conventional collective treatment also had a higher misuse than the two new criteria. In scenarios with a random pen allocation, the slope criterion was associated with the lowest misuse whereas in scenarios with sorted pens, the two new criteria were not significantly different.
Figure 6: Distributions of the proportion of treatment misuse. This proportion is the ratio of doses given to healthy individuals on the total AMU. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
The criterion for triggering a collective treatment affected the delay between the detection of the first BRD case and the collective treatment (Figure 7). Indeed, when collective treatments where triggered with the slope criterion, the delay between the first case and the collective treatment was shorter than a day in the majority of the stochastic replicates, whereas this delay was more often between one and two days for the conventional criterion. Collective treatment triggered by the severity criterion were triggered in less than a day in between 26% and 37% of the replicates with collective treatment. In around 43% of the cases, it was triggered between one and two days after the first case in the pen.
Figure 7: Delay between the first case in a pen and the collective treatment (in days). First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row: 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 90% of high risk animals. The effect of the treatment modality on the delay is significant according to a $\chi^{2}$ test.
On an individual level, collective treatments shortened the delay between the infection and the treatment (Figure8). The delay exceeded 2 days in less than 15% of the replicates in scenarios with collective treatments, while delays up to 9 days were observed in scenarios with individual treatments. Among collective treatment criteria, the conventional criterion yielded the biggest proportions of delays shorter than a day (up to 84% in HR90 scenarios), while the other two criteria yielded similar results.
Figure 8: Delay between infection and treatment (in days). First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row: the population is composed of 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: the population is composed of 10% of low risk animals and 90% of high risk animals.
@@
Pens containing animals of the same individual risk level were grouped together across the simulations to observe proportions of triggered collective treatment in each specific group (Table 3). A collective treatment was triggered in most of the pens composed of animals with high risk of respiratory disease, irrespective of the criterion. The conventional criterion had the highest proportion of triggered collective treatments in pens composed of medium risk individuals (83.9%), while the severity criterion triggered a collective treatment in 58.1% of these pens and the slope criterion only triggered a collective treatment in 33.8% of them. In pens composed of low risk individuals, collective treatments triggered by the conventional criterion happened in 35.9% of the pens, while collective treatments based on the severity and slope criteria only happened in 12.6 and 5.5% of the pens respectively.
@@
Table 3: Proportion of triggered collective treatments in sorted pens.
Sensitivity analysis
The between-pen contact rate had no impact on the outputs when it was lower than its nominal value (Figure 9). In scenarios with a random pen allocation, a contact rate equal to 0.1 was associated with significantly higher AMU and cumulative incidence only when the slope criterion was applied. However, the scenarios with sorted pens were more sensitive to changes in contact rate values: indeed, contact rates higher than the nominal value were associated with significantly higher AMU and cumulative incidences for every collective treatment stragies.
Figure 9: Analysis of the sensitivity of the outputs to between-pens contact rate values. Top panel: Cumulative incidence variation relative to the median of the baseline scenario – $c=0.01$ – (in %). Bottom panel: AMU variation relative to the median of the baseline scenario (in %). First row: the animals are allocated randomly in pens. Second row: the animals are allocated in pens according to their individual risk level. CT: Collective treatment. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction.
The transmission rate variations yielded a stronger impact on the cumulative incidence than on the AMU (Figure 10). Indeed, the AMU was significantly lower when the transmission rate was equal to 0.00375, with no other clear differences for the other transmission rate values. In particular, the AMU for the slope criterion remained unchanged for every transmission rate value. The cumulative incidence was similarly affected by lower transmission rate values. Additionnaly, it was also sensitive to transmission rates higher than the nominal value, especially in scenarios with sorted pens. Indeed, in scenarios with sorted pens, the cumulative incidence was significantly higher than the baseline scenario when the transmission rate was equal to 0.00625 for every collective treatment criterion.
Figure 10: Analysis of the sensitivity of the outputs to transmission rate values. Top panel: Cumulative incidence variation relative to the median of the baseline scenario – $\beta=0.005$ – (in %). Bottom panel: AMU variation relative to the median of the baseline scenario (in %). First row: the animals are allocated randomly in pens. Second row: the animals are allocated in pens according to their individual risk level. CT: Collective treatment. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction.
Changes in contact rates between pens had little to no impact on the ranking of the collective treatment criterion (Figure 11). Indeed, For every contact rate value, the slope criterion was always associated with the highest cumulative incidence in scenarios with random pens and was associated with the lowest AMU.
Figure 11: Impact of the collective treatment strategies on the outputs with the tested between-pen contact rate values. Top panel: Cumulative incidence. Bottom panel: AMU. First row: the animals are allocated randomly in pens. Second row: the animals are allocated in pens according to their individual risk level. Each column represents the tested value for the contact rate. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction..
Similar results were obtained when comparing the collective treatment criteria under several transmission rates (Figure 12). Indeed, the slope criterion was still associated with the lowest AMU and the highest incidence in random pen allocation scenarios.
Figure 12: Impact of the collective treatment strategies on the outputs with the tested transmission rate values. Top panel: Cumulative incidence. Bottom panel: AMU. First row: the animals are allocated randomly in pens. Second row: the animals are allocated in pens according to their individual risk level. Each column represents the tested value for the transmission rate. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction.
@@ Discussion
This study aimed at modeling the impact of triggering collective antimicrobial treatments according to three different criteria on the spread of M. haemolytica in multi-pens fattening systems while accounting for the multi-pen structure, individual risk of being affected by BRD and farming practices in terms of pens composition. We computed the cumulative incidence, the average duration of severe clinical signs and the mortality. We put these outputs in balance with the AMU and treatment precision for each scenario. We also computed the delay between the beginning of an episode and the beginning of the treatment and calculated the proportion of pens triggering collective treatments based on each criterion. We proposed that a novel criterion, based on the speed of the spread rather than on a proportion of detected individuals, could yield a better trade-off between BRD case incidence and AMU than the conventional way of implementing collective treatment. This result could have interesting operational implications on the decision-making regarding collective treatment in fattening farms.
@@
Our results showcased a positive impact of collective treatments on disease related mortality in the herd in high risk scenarios. However, no criterion was identified as systematically better than the others in reducing the mortality. This can be due to the fact that collective treatments happened rather rapidly after an outbreak, in spite of some differences in delay. This swift reaction did not let time for animals to develop severe clinical signs and to die subsequently to them. Our results aligned with field studies showingreduced mortality in herds where collective treatment was used ([oconnor_mixed_2013]).
@@
The supplementary cost in antimicrobial was less desirable when the general risk was lower in the population, as the reduction in incidence and severity thanks to collective treatment was less substantial. The overuse of antimicrobial leads to antimicrobial resistance and treatment purchase and administration are costs supported by the farmer ([laxminarayan_antibiotic_2013]). Therefore, there appears to be a trade-off between incidence reduction and antimicrobial use which could benefit from a characterization of the risk level of the pens. Such a characterization could guide the farmers in making the best decision regarding treatment and herd management ([edwards_control_2010]).
@@
Our model allowed us to identify a beneficial novel criterion for collective treatment. Indeed, instead of a threshold of detected animals in a pen, basing a collective treatment decision on the speed of the spread of the disease allowed marginal gains in disease incidence while drastically reducing the AMU and the treatment misuse when compared to the other collective treatment criteria. A field meta-analysis of randomized clinical trials showed that the risk reduction of BRD varied drastically with the criterion for collective treatment ([baptiste_antimicrobial_2017]). The routine criterion is 10% of morbidity for two consecutive days, however, there is no evidence based studies proving the relevance of this threshold ([radostits_herd_2001]). Perspectives of this study include the investigation of the novel treatment strategies in broader contexts in order to translate the thresholds used into operative indicators.
@@
Collective treatments based on the speed of the spread of the pathogen proved to be triggered sooner than with the other criteria. This precocity was key for the reduction of the incidence and severity of the disease, especially in high-risk populations. It also showed to target more specifically high-risk groups than the other criteria, as it reduced the proportion of low-risk pens having received collective treatments. On an individual level, the delay between the infection and the administration of the first dose of treatment was also characterized. This delay was shortened by the implementation of collective treatment strategies, particularly for the conventional criterion. These are original findings with clear practical implications, especially with regards to the need for a better knowledge by farmers of individual risk of animals developing BRD, in order to account for this information when constituting pens. However, triggering collective treatments greatly relies on the ability to detect diseased animals in the first place, therefore, increasing the sensitivity via standardized detection procedures may increase the benefit of the subsequent control measures and may modify the most convenient criteria to trigger collective treatment ([kamel2024strategies]).
The novel slope criterion proposed by this study also outperformed the conventional criterion in terms of antimicrobial misuse, especially in moderate risk scenarios. As the animals with a high individual risk of BRD are more prone to develop and to transmit respiratory pathogens, targetting the most at risk pens is a desirable trait for BRD control measures. Our results on collective treatments occurrence in sorted pens highlight the ability of the novel slope criteria to target the most at risk pens, therefore reducing the antimicrobial use in pens with a lesser risk of an outbreak. This also translated to the lowest proportion of treatment misuse at herd level when compared to the other two collective treatment criteria. In scenarios with individual treatments only, the misuse of antimicrobial on healthy animals was even lower than for the slope criterion, as it was solely due to the false discovery of diseased animals via temperature check, as hyperthermia could be non specific to infection.
The conclusions brought up by our model proved to be robust to changes in transmission rates and between-pen contact rates. Furthermore, it also brought evidence that the slope criterion benefited from sorted pens, as this setting allowed the incidence and AMU to be particularly stable. These findings allowed us to prove the robustness of our results and to further emphasize the potential of the new collective treatment criterion we brought forward.
@@
There is clear evidence that antimicrobial resistance is rising among M.haemolytica populations ([woolums_multidrug_2018, timsit_prevalence_2017]). This rise happens on time scales that are longer than the one we modeled. Therefore, we modeled the existence of resistant strains by calculating a probability of treatment success, which value was set along the span of the simulation. Future work could aim at evaluating the long term effects of collective treatment protocols on the antimicrobial trends.
The force of infection used in this model assumed an increased pathogen transmission within pens with an additional equal contribution of the surrounding pens. This hypothesesis relied on the fact that aerosolization of M.haemolytica has been pointed as a contributing factor to rapid disease spread among cattle populations in close quarters ([rice_mannheimia_2007]). However, this airborne transmission is not as effective as transmission via close contacts. Adaptions of the model could therefore implement a force of infection based on the farm spatial organization ([ackermann2000response]).
Our study advocates for more transparency on the individual risk level for BRD in the cattle production system. Such a characterization could allow mindful investments for the farmers as well as more adequate control measures regarding housing and treatment. Indeed, our novel collective treatment criterion proved to be able to target at-risk pens while not applying collective treatments to lower risk pens.
@@ Acknowledgements
Pauline Ezanno is recommender of PCI Animal Science.
@@
The complete code for the model model can be found on this git repository: https://archive.softwareheritage.org/browse/origin/directory/?branch=refs/heads/PCI_animal_science2024&origin_url=https://forgemia.inra.fr/dynamo/brd-models/brd-public.git. The version of the model used for this work is tagged under pci_animal_science_2024.
Annexe A SI1: Model processes
A.1 Implementation and formalism of the model
@@
The model is a stochastic individual based model. In the model, five processes drive the states of individuals: infection, hyperthermia, clinical signs, detection and treatment. They are represented by a formalism broadly used in computer science, finite state machines, which is close to flow diagrams used by epidemiologists, with a higher expressiveness. They represent finite state an individual can find itself in. Their textual description in the YAML format in the model file is easy to read and can be automatically converted in a graphical representation by EMULSION. The model can be found on a git repository, accessible by following this link: https://forgemia.inra.fr/spicault/brd-withinherd-emulsion.git. The version of the model used for this work is tagged under pci_animal_science_2024.
@@
To help understand the state machine diagrams, we present how the diagram summarising the infection process relates to the state machine specification, extracted from the whole model. More details on how state machines differ from flow diagrams are available in the documentation: https://sourcesup.renater.fr/www/emulsion-public/pages/Modelling_principles.html.
Annexe B SI2: Threshold selection for the Novel criteria for collective treatment
B.1 Method